OpenStudy (anonymous):

Multiply. (2u-6x)(4u-3x+2)

OpenStudy (whpalmer4):

Multiply the first term of (2u-6x) by each term in (4u-3x+2). Then multiply the second term of (2u-6x) by each term in (4u-3x+2). Add up all the pieces and simplify by collecting like terms. In general, you multiply each term of each multiplier by each term of the other. \[(a+b+c)(d+e+f) = a(d+e+f) + b(d+e+f) + c(d+e+f) = \]\[ad + ae + af + bd + be + bf + cd + ce + cf\]

OpenStudy (anonymous):

so like.... 2u(4u-3x+2) - 6x(4u-3x+2)

OpenStudy (whpalmer4):

Exactly. Now expand those two products, and you get?

OpenStudy (anonymous):

so its 12u^2-30ux+18x-12

OpenStudy (whpalmer4):

Not quite. \[2u(4u-3x+2) = 2*4*u*u-2*3*u*x+2*2*u = 8u^2-6ux+4u\] \[-6x(4u-3x+2) = -6*4*x*u -6(-3)*x*x -6*2*x = -24ux + 18x^2 -12x\] Adding them together, \[8u^2-6ux+4u-24ux+18x^2-12x = \]

OpenStudy (anonymous):

so, then I add like terms

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

Ok! thanks for your help

OpenStudy (whpalmer4):

What do you get for your result? you can just tell me which terms combine and what they combine to...

OpenStudy (anonymous):

I got 12u^3-30ux^2+18^2-12x

OpenStudy (whpalmer4):

How did you get a u^3 term?

OpenStudy (anonymous):

8u squared and 4u

OpenStudy (whpalmer4):

Look at my unsimplified result: \[8u^2-6ux+4u-24ux+18x^2-12x\]The only two terms which can be combined are -6ux and -24ux. You can't add 8u^2 + 4u like that.

OpenStudy (whpalmer4):

You can only add them together if they have the same base and exponent.

OpenStudy (whpalmer4):

8u + 4u = 12u 8u^2 + 4u^2 = 12u^2 8u^2 + 4u = 8u^2 + 4u (or you could write it as 4u(2u+1))

OpenStudy (anonymous):

oh, i thought they were like terms as long as they had the same variable

OpenStudy (whpalmer4):

As an example, suppose u = 3 8u^2 = 8*3*3 = 72 4u = 4*3 = 12 8u^2 + 4u = 72 + 12 = 84 However, with the statement that 8u^2 + 4u = 12u^3, you are saying that 72 + 12 = 12*3*3*3 = 108

OpenStudy (whpalmer4):

Nope, have to have both the same base and the same exponent.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

With that in mind, what is your new, improved answer?

OpenStudy (anonymous):

8u^2-30ux^2+4u+18x^2-12x

OpenStudy (whpalmer4):

Now where did you get -30ux^2 from? Details are very important in algebra...

OpenStudy (anonymous):

-6ux and -24ux

OpenStudy (whpalmer4):

Add them together and the exponents do not change. -6ux -24ux = -30ux.

OpenStudy (anonymous):

so its just -30ux

OpenStudy (whpalmer4):

Yes. Final answer is \[8u^2-30ux+4u+18x^2-12x\]

OpenStudy (anonymous):

that was really long

OpenStudy (whpalmer4):

If we knew a value in place of x or u, it would collapse down to something smaller. But we don't...

OpenStudy (anonymous):

ok. do you mind helping me with another problem?

OpenStudy (whpalmer4):

Post it as another problem, I have to leave...

OpenStudy (anonymous):

ok. cool